OpenStudy (anonymous):

One number is 4 more than another, and their sum is 60. What is the smaller number? 26 28 32

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

28?

Vocaloid (vocaloid):

good

OpenStudy (anonymous):

thanks:)

OpenStudy (trojanpoem):

x + y = 60 x - y = 4 2x = 64 x = 32 y = 60 -32 = 28